OpenStudy (anonymous):

PLEASE HELP I AM STUCK !!!!!!!!!!!!!! Which of the following equations best represents the graph of f(n) If f(x) = 7x - 1, what does f(12) represent

jimthompson5910 (jim_thompson5910):

To find f(12), replace every x with 12 and evaluate.

OpenStudy (anonymous):

i dont get it.

jimthompson5910 (jim_thompson5910):

f(x) = 7x - 1 f(12) = 7*12 - 1 ... replace every x with 12 now evaluate

OpenStudy (anonymous):

f(x)=83?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

f(12) = 83 actually

jimthompson5910 (jim_thompson5910):

but you got it pretty much

OpenStudy (anonymous):

oh ok thank you. well you help me on others ?

jimthompson5910 (jim_thompson5910):

I can help with one more

OpenStudy (anonymous):

The first five terms of a sequence are shown below: 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence?

jimthompson5910 (jim_thompson5910):

what do you notice about the jump from 5 to 9? From 9 to 13? etc etc

OpenStudy (anonymous):

a. f(n) = 5n - 1; n ≥ 1 b. f(n) = 5n + 1; n ≥ 1 c.f(n) = 4n - 1; n ≥ 1 d. f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

it goes by 4

jimthompson5910 (jim_thompson5910):

so that's the common difference

jimthompson5910 (jim_thompson5910):

which means d = 4

jimthompson5910 (jim_thompson5910):

the starting term is 5, so a1 = 5

jimthompson5910 (jim_thompson5910):

nth term: an = a1 + d(n-1) an = 5 + 4(n-1) an = 5 + 4n-4 an = 4n + 1 this is where n is an integer and it starts at n = 1

jimthompson5910 (jim_thompson5910):

The formula an = a1 + d(n-1) is the general arithmetic sequence (the nth term of it)

jimthompson5910 (jim_thompson5910):

a1 is the first term d is the common difference (telling you how much to increase each term each time)

OpenStudy (anonymous):

so it would be f(n)=4+1 n_>1

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you and for the first question, its not what does f(12) euqals its what it represents.....

jimthompson5910 (jim_thompson5910):

it's the y value or output of the function when the input is x = 12

OpenStudy (anonymous):

a. The value of x when (7x - 1) = 12 b. The value of (7x - 1) when x = 12 c.The value of 12 f(x) when x = 12 d.The value of f(x) when (7x - 1) = 12 i think its b

jimthompson5910 (jim_thompson5910):

in goes x = 12, out comes y = 83

jimthompson5910 (jim_thompson5910):

b is correct

OpenStudy (anonymous):

thank you so much ! its clear now.

jimthompson5910 (jim_thompson5910):

you're welcome